### v2

# AI Research Agents for Machine Learning: Search, Exploration, and Generalization in MLE-bench

Abstract
AI research agents are demonstrating great potential to accelerate scientific progress by automating the design, implementation, and training of machine learning models. We focus on methods for improving agents’ performance on MLE-bench, a challenging benchmark where agents compete in Kaggle competitions to solve real-world machine learning problems. We formalize AI research agents as search policies that navigate a space of candidate solutions, iteratively modifying them using operators. By designing and systematically varying different operator sets and search policies (Greedy, MCTS, Evolutionary), we show that their interplay is critical for achieving high performance. Our best pairing of search strategy and operator set achieves a state-of-the-art result on MLE-bench lite, increasing the success rate of achieving a Kaggle medal from 39.6 % to 47.7 %. Our investigation underscores the importance of jointly considering the search strategy, operator design, and evaluation methodology in advancing automated machine learning.
[…]
1 Introduction
Figure 1: AIRA agents use the AIRA-dojo environment, AIRA operators, and search policies to achieve state-of-the-art performance on MLE-Bench lite. Operating in AIRA-dojo improves the performance of AIDE222—the previous state-of-the-art from Chan et al. (2025). Enhanced operators lead to an improvement from AIDEgreedy to AIRAgreedy. Exploring the space of search policies can yield further performance gains, only when these constraints are addressed.
22footnotetext: o1-preview is deprecated.
Science is based on searching the open-ended space of hypotheses and testing them in a controlled experiment (Langley et al., 1987). Recent breakthroughs have resulted in artificial intelligence (AI) agents that offer great potential to automate the scientific discovery process (Yamada et al., 2025; Swanson et al., 2024; Boiko et al., 2023). A key obstacle to improving research agents is that their designs entangle several factors for performance, making it difficult to pinpoint sources of improvement via controlled experiments at scale. These factors span algorithm design, concrete implementation, and ability to leverage compute, as performance gains accrue only if no layer in the stack bottlenecks the benefits from additional compute resources (Fig.˜7). This challenge is exemplified in MLE-bench (Chan et al., 2025), a benchmark where AI agents compete in Kaggle competitions to solve real-world machine learning (ML) problems. Notably, the state-of-the-art approach AIDE (Jiang et al., 2025) does not fully disentangle these performance factors, providing limited insight into which components primarily drive the agent’s performance and where improvements are needed.
Figure 2: Overview of AIRA. Given a problem specification, AIRA maintains a search graph whose nodes are (partial) solutions. At each iteration, the agent (1) selects nodes via a selection policy, (2) picks an operator via an operator policy and applies this operator to the node, and (3) scores the resulting solution via a fitness function. Here, a greedy node‐selection strategy applies the improve operator to the highest‐scoring node.
We start by formalizing the design of AI research agents as a search algorithm composed of two components: The first one is the search policy which is used to navigate the space of candidate solutions, and the second is the operators which iteratively modifies existing solutions to generate new candidate solutions. In this framework (Fig.˜2), AIDE is represented as a greedy search algorithm that, at each step, applies one of its code operators (i.e., Draft, Debug, Improve) to the current best solution. This allows us to disentangle the effect of the search from that of the operators.
To assess alternative search algorithms, we develop more sophisticated agents performing Evolutionary and Monte Carlo Tree Search (MCTS) (De Jong, 2017; Kocsis and Szepesvári, 2006). We empirically show that AIDE’s operators, rather than the search algorithm, are a bottleneck to better performance.
Based on these findings, we design an improved set of operators and evaluate them when paired with the above-mentioned search algorithms. Our best-performing agent achieves state-of-the-art results on MLE-bench lite, increasing the rate of achieving a Kaggle medal from 39.6 % to 47.7 %.
Additionally, we investigate how the generalization gap—the difference between validation and test scores—affects the agents’ performance. In ML engineering tasks, as in many real-world use cases, an agent can access only proxy metrics (e.g. validation loss) to guide the search process, while the final solution is evaluated on a held-out test set. Therefore, a gap between the expected (validation) and actual (test) loss, could potentially mislead the search process. Indeed, we find systematic overfitting: selecting the final solution in a search graph by its test—rather than validation—score, would increase the medal rate by 9 to 13 % (absolute scale), depending on the search algorithm. These findings highlight the importance of rigorous evaluation protocols during search and regularization for robust and scalable scientific discovery.
Finally, to conduct experiments, we develop AI Research Agent dojo (AIRA-dojo), a framework that provides a scalable and customizable environment for AI research agents. First, AIRA-dojo exposes a robust and flexible interface to compute resources, which is essential for building effective agents. The baseline, AIDE, implemented in AIRA-dojo achieves a performance increase of 10.68 % (absolute scale) over the reported results (Chan et al., 2025). Second, AIRA-dojo enables users to experiment with custom operators, search policies, evaluation methods, and tasks within a comparable setup. This facilitates a rigorous scientific study of AI research automation.
Rerunning the best-performing agents with our latest version of AIRA-dojo further increased performance to 55 % on MLE-Bench Lite. For details on these experiments and results on the complete MLE-bench suite of tasks, see Section˜11.
2 Research Agents as Search Algorithms
AI research agents typically approach machine learning problems by generating artifacts – codebases designed to solve a given task. Executing these artifacts yields trained models, which are then evaluated against a chosen performance metric. The agent iteratively refines its artifacts to improve performance on this metric.
Previous research indicates that LLMs alone are insufficient to effectively solve such open-ended tasks (Nathani et al., 2025). In particular, LLM performance significantly improves when augmented with external tools (Qin et al., 2024), execution feedback (Gehring et al., 2025), and solutions addressing context limitations. Therefore, developing high-quality models typically involves iterative experimentation, where insights from prior experiments inform subsequent refinements. Recent advancements by Jiang et al. (2025) demonstrate state-of-the-art performance by conceptualizing this iterative experimentation as a tree search over potential solutions.
In this section, we formalize and generalize this perspective by modeling research agents as graph-based search algorithms. The proposed framework allows systematic exploration of alternative agent design choices, providing insights into how different algorithms affect the exploration-exploitation trade-off—a fundamental aspect of search (Kocsis and Szepesvári, 2006; Črepinšek et al., 2013).
2.1 Graph–based Search Framework
We consider an agent that operates by searching a directed graph $\mathcal{G}_{t}=(V_{t},E_{t})$ that evolves over multiple iterations $t=0,1,\dots$, where each node $v\!\in\!V_{t}\subseteq\mathcal{S}$ represents an artifact belonging to the set of all possible artifacts $\mathcal{S}$, while each directed edge $(v_{i},v_{j})\!\in\!E_{t}$ represents a transformation from $v_{i}$ to $v_{j}$. The root $v_{0}$ is the initial artifact that can represent an empty or starting artifact.
Definition 1 (Components of the Search Algorithm). A graph‑based search algorithm is specified by the tuple $\bigl(\mathcal{F},\pi_{\text{sel}},\mathcal{O},\pi_{\text{op}},\tau\bigr)$:
  • Fitness Function. $\mathcal{F}:\mathcal{S}\to[0;1]$ is a function that estimates the value or quality of a node $v\in V_{t}$. Since true fitness is typically not available, $\mathcal{F}$ is often a proxy measure of the value of the node.
  • Selection Policy. $\pi_{\text{sel}}:2^{V_{t}}\to\!2^{V_{t}}$ chooses a subset of nodes $U_{t}\subseteq V_{t}$ on which to operate, typically guided by a heuristic function $h:V_{t}\rightarrow\mathbb{R}$, which assigns a scalar estimate to each node. The heuristic may be derived directly from the fitness function $\mathcal{F}$, such as the upper confidence bounds for trees (UCT) used in MCTS (Kocsis and Szepesvári, 2006), or other custom heuristics tailored to the domain (Crawford and Auton, 1996; Hooker and Vinay, 1995).
  • Operator Set. $\mathcal{O}=\{\,o_{\ell}:2^{\mathcal{S}}\to\mathcal{S}\}_{\ell=1}^{L}$ comprises $L$ transformation functions that propose new artifacts $v=o_{\ell}(\{v_{k}\}_{k=1}^{m})\in\mathcal{S}$ from one or more selected artifacts. In AIDE, examples include draft, debug, and improve instantiated as prompt-based instructions to an LLM. Composite operators (e.g. the result of applying a sequence of base operators) can also be used.
  • Operator Policy. $\pi_{\text{op}}:\mathcal{O}\times U_{t}\to\mathcal{O}$ decides which operator to apply to the current node selection.
  • Termination Rule. $\tau$ halts the search when a computational budget is exhausted, progress stalls, or a fitness threshold is reached.
At each iteration, the agent selects existing promising artifacts, applies transformation operators, and updates the search graph, propelling the discovery process forward. We discuss specific instantiations of search algorithms in Sections 2.3, 4.2. In all our instantiations: 1 the termination criterion is set as the wall-clock time or the maximum number of artifacts, whichever happens first; 2 the fitness function $\mathcal{F}$ is defined for each node by the operator that generates or modifies it (i.e., $\mathcal{F}$ is not global); 3 all operators are LLM-driven except for the Memory operator, which is defined by hand.
2.2 Operators
We define operators as high-level functions that take in existing artifacts and produce new ones. These can range from simple rule-based parsers to LLM calls using prompting techniques (Wei et al., 2022; Yao et al., 2023b) or more complex agents like Cursor (Inc., 2025).
This broad definition enables a unified comparison across search methods. A search algorithm can be instantiated with any mix of LLM-based, tool-based, or nested search operators.
Building on the demonstrated effectiveness of AIDE (Jiang et al., 2025), we adopt a similar operator set as introduced in their framework, consisting of the following: 1 Draftinitializes the search process by generating an initial population of candidate artifacts. 2 Debugattempts to identify and correct errors in invalid artifacts. 3 Improverefines valid artifacts to enhance their performance according to the evaluation criteria. 4 Memorychooses how and where information from past artifacts is used in subsequent operations. 5 Furthermore, we introduce Crossover which recombines useful elements from two artifacts to create a new candidate.
Section˜3 contains further implementation details.
2.3 Re-casting AIDE in our Notation
AIDE (Jiang et al., 2025) is an LLM-driven agent that frames problem-solving as a tree-search over Python scripts. In the tuple $(\mathcal{F},\pi_{\text{sel}},\mathcal{O},\pi_{\text{op}},\tau)$ from Section˜2.1 its components are:
Fitness & selection ($\mathcal{F},\pi_{\text{sel}}$). For each node $v$, fitness is the mean 5-fold cross-validation (CV) score $\mathcal{F}(v)\!\in\![0,1]$ [^3] In practice, CV scores are not necessarily bounded between zero and one; for example, RMSE is an unbounded metric.. The selection policy is greedy with respect to $\mathcal{F}$. This means at iteration $t$ the agent selects and operates on
$v^{\star}=\arg\max_{v\in V_{t}}\mathcal{F}(v),$
[…]
Operator set $\mathcal{O}_{\text{AIDE}}$. The agent exposes three LLM operators {Draft, Improve, Debug} and one handcrafted operator Memory or as defined by Jiang et al. (2025)—the Summarization operator. The operators are designed to output the following: Draft (3–5-sentence plan + fenced script that trains, evaluates, and writes submission.csv); Debug (short diagnosis and repaired script given a traceback); Improve (Creating exactly one measurable change — feature, architecture, schedule, etc. — in plan + code form); Memory (running summary of all previous designs, scores, and notes that is appended to every Draft/Improve prompt).
Operator policy $\pi_{\text{op}}$. (a) Seeding: invoke Draft exactly $n_{d}$ times from the root $v_{0}$; (b) Logging: after every Draft or Improve, call Memory; (c) Main loop: for the node chosen by $\pi_{\text{sel}}$ apply Improve if it is valid and at least $n_{d}$ drafts exist, if less than $n_{d}$ draft nodes exist, Draft is called, otherwise apply Debug.
[…]
3 Experiment Design
3.1 AIRA-dojo
An agent’s environment greatly influences its performance. To systematically study the space of agentic policies (see Fig.˜7), we introduce the AI Research Agent (AIRA) dojo—a scalable and customizable framework for AI research agents. AIRA-dojo provides the environment in which agents operate, along with abstractions for operators and policies as described in Section˜2, and tasks that define evaluation criteria for agent performance. Using these abstractions, we implement and evaluate four search policies: MCTS, Evolutionary, Greedy, and our own implementation of AIDE. We hope that AIRA-dojo’s scalability and customizability, together with the provided agent and task implementations, will support and advance future research in the community.
The infrastructure design of AIRA-dojo was informed by several reliability and performance constraints, as discussed in more detail in Section˜14.
3.2 Environment
The environment defines the context in which agents operate.
Action Space. We use Jupyter notebooks to execute code from the agents. With this interface, agents can execute arbitrary Python code, perform file reads and writes, and even use the shell via Jupyter magic commands. The environment captures and returns the status, standard outputs, tracebacks for debugging, and the code block execution time to the agents.
Isolation. The environment enforces program-level constraints, such as total runtime limit, GPU and CPU usage, memory, and storage, using Apptainer containers (Kurtzer et al., 2021). This ensures complete isolation from host systems, preventing agents from affecting the host environment or other agents, and mitigates risks from unintended actions or failures, such as data leakage or interference with other processes. Furthermore, agents possess root-like privileges within their isolated containers, granting them full control over their environment. This allows them not only to configure their environments using standard package management tools such as pip, conda, but even apt-get install.
Superimage. Superimage, a container image, provides a base set of tools required for common ML tasks. This image includes pre-configured CUDA support, key deep learning frameworks such as PyTorch and TensorFlow, and essential data science libraries. Each coding session starts from the original Superimage state and can diverge based on the agent’s actions while not affecting the state of the other agents.
Together, these design choices create a stable and robust testbed that eliminates confounders at both the system and implementation levels. This reliability enables reproducible benchmarking and facilitates the development of long-running agentic systems across thousands of parallel runs and diverse tasks.
3.3 MLE-bench
MLE-bench (Chan et al., 2025) comprises of 75 tasks sourced from Kaggle, designed to assess the capabilities of agents in machine learning engineering. In our experiments, we evaluate on MLE-bench lite—a curated subset of 22 tasks selected from the full benchmark.
Our experiments showed that existing methods exhibit high variance in this benchmark (see Section˜16). Focusing on a smaller subset of tasks allows us to allocate more seeds per task and increase our confidence in our results.
In line with the benchmark guidelines, we assess each agent’s performance using the Medal Success Rate. Specifically, for each task, agents earn a bronze, silver, or gold medal according to task-specific percentile thresholds. We report the percentage of attempts in which an agent secures a medal.
3.4 Experimental Details
Environment. Each candidate agent is launched in a freshly initialized, sandboxed process in AIRA-dojo. This guarantees that file systems and environment variables are isolated across evaluations, and there is no cross-agent interference. Every sandbox is provisioned with a fixed hardware quota: 1 dedicated H200 GPU, 24 logical CPU cores, 100 GB of RAM, and 1 TB of additional scratch storage. Internet access is permitted solely for fetching third‑party packages and model checkpoints.
Time Constraints. In line with MLE-bench, the agent is allowed a 24-hour wall-clock window. Within this period, each agent has a maximum runtime of 4 hours per code execution. We chose to reduce this from the 9-hour limit used in MLE-bench after preliminary experiments showed no difference in performance and a higher average number of valid nodes in the search trees.
LLMs. We conducted all the experiments with the full-sized DeepSeek R1 (DeepSeek-AI, 2025) model, with 128K-token context window to ensure input coverage without truncation. Due to wall-clock constraints, this choice was guided by both the model’s capabilities and applicable rate limits. In particular, we selected DeepSeek R1 as the most capable open-source model available, which allowed us to self-host inference servers and maintain high experimental throughput without encountering rate limits. For the main results, we also evaluated o3 (OpenAI, 2024), one of the most capable closed-source models. To ensure experiment validity and avoid hitting rate limits, we limited the number of parallel runs when experimenting with o3. We always use GPT-4o (OpenAI, 2024) with Structured Outputs to parse code execution outputs—extracting run success, summary text, and validation metric. For the figures, we take [OAI] AIDE o1 artifacts from the MLE-bench Chan et al. (2025) GitHub repository. To enable a direct comparison with their results, we also evaluated o1-preview. However, we were only able to complete experiments for one agent before the model was deprecated.
See Section˜12 for further implementation details.
4 AIRA
4.1 Operators $\mathcal{O}_{\text{AIRA}}$
As part of AIRA-dojo, we propose a new operator set based on $\mathcal{O}_{\text{AIDE}}$, which we refer to as $\mathcal{O}_{\text{AIRA}}$. In designing $\mathcal{O}_{\text{AIRA}}$, we focus on maintaining a cleaner context through better-scoped memory, and encouraging structured reasoning and strategic diversity in ideation. The key differences between the two sets of operators are:
  1. 1
    Prompt‑adaptive complexity. We introduce a dynamic complexity cue within the system prompt in order to guide the complexity of artifacts generated by the draft and improve operators. The complexity is determined by the number of children, $n_{\mathrm{c}}$, of the node being processed:
    $\mathrm{complexity}(n_{\mathrm{c}})=\text{``minimal''}\penalty 10000\ \mathrm{if}\penalty 10000\ n_{\mathrm{c}}<2,\penalty 10000\ \text{``moderate''}\penalty 10000\ \mathrm{if}\penalty 10000\ 2\leq n_{\mathrm{c}}<4,\penalty 10000\ \text{``advanced''}\penalty 10000\ \mathrm{if}\penalty 10000\ n_{\mathrm{c}}\geq 5$
    For the draft operator, this cue influences the complexity of the generated ideas. For the improve operator, it guides the complexity of the enhancements. This dynamic signal helps prevent premature over-engineering by ensuring the agent provides simple solutions when appropriate, while encouraging more thorough exploration when more advanced solutions may be necessary.
  2. 2
    Scoped memory. We modify the memory operator to extract different types of memories depending on the operator used. Specifically, for draft and improve, it retrieves only sibling memories—the children of the artifact the agent is applying the operator to—thereby promoting diversity. This scoped memory approach prevents overloading the context and reduces behavior indicative of mode collapse. Conversely, for debug, it retrieves the entire ancestral memory of the artifact’s debug chain, enabling review of prior fix attempts and avoiding “undo–redo” oscillations.
  3. 3
    Think Tokens. For reasoning models, we use the operators’ system prompts to explicitly encourage them to use thinking tokens for structured reasoning and reflection. These thoughts are stripped from the final answer—remaining invisible to other operators (such as memory). On average, we observe a $2\times$ increase in completion tokens generated by the AIRA operator set (see Section˜13).
4.2 Agents
In this section, we introduce three agents that combine the operators $\mathcal{O}_{\text{AIRA}}$—proposed in Section˜4.1—with distinct search policies (see Appendix 15 for the rationale behind their selection). Each agent uses the same proxy–fitness function $\mathcal{F}$ (5-fold CV) and the same termination criterion $\tau$ (based on wall-clock time or artifact cap).
AIRAgreedy. This agent employs greedy search (Section˜2.3) using the $\mathcal{O}_{\text{AIRA}}$ operator set. Any performance improvement of AIRAgreedy over AIDEgreedy directly reflects the benefit of the new operators, since the only difference between the two is the operator set.
AIRAmcts. This agent uses Monte-Carlo Tree Search (MCTS) Kocsis and Szepesvári (2006) with $\mathcal{O}_{\text{AIRA}}$ operator set. Our implementation of MCTS follows the canonical loop (selection, expansion, evaluation, and backup), but omits simulated roll‑outs: the leaf value of expanded nodes is given directly by the proxy fitness function $\mathcal{F}$:
  • Selection. From the root node $v_{0}$, descend by selecting the node with the highest UCT score $\pi_{\text{sel}}(v)=\arg\max_{v\in V_{t}}h_{\text{UCT}}(v)$ where $h_{\text{UCT}}(v\!\mid u)=Q(v)+c\sqrt{\log N(u)/(N(v)+\varepsilon)}$, where $N(\cdot)$ and $Q(\cdot)$ are the visit count and running mean fitness. Here, $u$ is the parent of $v$.
  • Expansion. Only leaves are expanded. The chosen operator from $\mathcal{O}_{\text{AIRA}}$ is applied $n$ times, creating $n$ children. Buggy children enter an automatic Debug loop until fixed or the budget expires.
  • Evaluation & backup. The leaf fitness is $\mathcal{F}(v_{\ell})$ is back-propagated to ancestors with the standard incremental update of $(N,Q)$.
AIRAevo. The evolutionary agent keeps a population $V_{t}$ of fixed size $n$ and repeats:
  • Parent selection: Select individuals with probability $\pi_{\text{sel}}(v)=\mathcal{F}(v)/\sum_{u\in V_{t}}\mathcal{F}(u)$.
  • Reproduction: With a fixed probability, apply Improve; otherwise, apply Crossover. Parents that are buggy are first processed by Debug until they are either fixed or the debug attempt limit is reached.
  • Replacement: Offspring replace the least-fit individuals in $V_{t}$.
For both AIRAmcts and AIRAevo, we normalize all fitness values using the minimum and maximum values observed during the search process. This normalization ensures a consistent set of hyperparameters throughout the search. To select the final solution, the AIRA and AIDE agents return the one with the highest validation score.
5 Experiments and Results
Figure 3: Searching with AIDE’s operators. When limited to AIDE’s operator set $\mathcal{O}_{\text{AIDE}}$, agents using more advances search policies (e.g., MCTS, evolutionary algorithms) gain no advantage, underscoring the operator set as the bottleneck.
5.1 Analyzing the Performance of the Current SoTA
The most effective search algorithms strike a balance between exploration and exploitation. In this section, we analyze the exploration–exploitation trade-offs of AIDE, the current state-of-the-art method, and then investigate how additional computation time increases its performance.
The three main factors that influence this balance are memory, the operator set, and the selection policy. Memory structures prior knowledge—storing promising solutions and tracking which regions of the solution space have been sampled—to inform subsequent decisions. This information can bias the search toward exploration by encouraging diversity or toward exploitation by discouraging it. Operators then apply controlled transformations to existing solutions: for instance, random mutations introduce diversity, targeted refinements exploit known strengths, and recombinations merge features from multiple candidates. Finally, the selection policy balances exploiting high-quality regions of the solution space with exploring less-tested areas by determining how to allocate computational resources. A non-greedy selection policy, such as in MCTS, periodically directs resources toward branches that may appear suboptimal, with the goal of uncovering (and subsequently exploiting) better solutions.
What is the effect of memory? To quantify the impact of the memory operator, we conduct a controlled ablation comparing the performance of AIDE with and without the memory operator enabled. As shown in Fig.˜3, both variants achieve nearly identical mean medal rates. This suggests that memory is not a driving factor behind AIDE’s strong performance.
[…]
Figure 4: a) AIDE’s performance profile over 24-hour search window. Perceived vs. actual medal rate over 24 hours of AIDEgreedy. The curves show the mean validation (agent-reported) and held-out test medal rates across 20 seeds for all tasks. The widening band illustrates the generalization gap, revealing how apparent gains on the validation set can mask overfitting and ultimately undermine the search process. b) Performance profiles of all agents after 24-hour search window.
Figure 5: Medal rates on MLE-bench Lite. Performance is shown for three medal categories: any medal, silver medals and above, and gold medals only. Error bars represent 95% confidence intervals computed using stratified bootstrapping.
What is the effect of exploration at the search level with AIDE’s operators? AIDE uses a selection policy that does not explore and always greedily selects the most promising candidate. To evaluate the impact of the search policy as a modulator of the exploration-exploitation tradeoff when searching using AIDE’s operators, we replaced the operators in AIRAmcts with that of AIDE and varied the UCT exploration constant $c_{\mathrm{UCT}}\in\{0,0.25,0.75\}$. The $c_{\mathrm{UCT}}$, which modulates the degree of exploration in MCTS, allows us to isolate the effect of search-level exploration on downstream performance. The resulting medal rates (Fig.˜3) showed only marginal differences across all $c_{\mathrm{UCT}}$ values, supporting our hypothesis that search-level exploration is constrained by the interaction between the operator set and the search policy. We observe the same limitation when performing the same operator replacement in AIRAevo.
What is the performance profile of AIDE? To examine the improvement gains over time, we plot AIDE’s anytime performance, which is the average medal rate achieved by the agent if the search was to terminate at time $t$. The results, summarized in Fig.˜4, suggest that while the agent’s perceived performance (validation score) continues to improve over time, the true test performance plateaus, even slightly decreases over time. These findings suggest that overfitting might be a fundamental limitation to the agent’s performance. We further investigate this in Section˜5.3.
5.2 AIRA Beyond Greedy
Based on the observation that search policies yield no performance gain with AIDE’s operators, this section is divided into two parts: 1 assessing the effectiveness of our improved operators (see Section˜4.1), 2 examining their interplay with the more advanced search policies. The results are summarized in Fig.˜5, with a detailed breakdown of performance for each task in Section˜18.
Evaluating the environment. But first, we highlight the benefits from AIRA-dojo. Specifically, our baseline agent implementation, AIDEgreedy o1-preview, operating in the AIRA-dojo environment, improves the medal rate from 35.2 % to 45.9 % compared to the reported results (Chan et al., 2025). This corresponds to state-of-the-art performance with a relative improvement of 30 % in medal rate. It is notable that AIDEgreedy o1-preview, outperforms both AIDEgreedy R1 and AIDEgreedy o3, despite o3 being the newest model in the series. While evaluating all agents with o1-preview would have been informative, we were only able to complete the AIDEgreedy experiment before the model was discontinued.
Evaluating the operator sets. AIDEgreedy and AIRAgreedy employ the same search policy but differ in their operator sets. Comparing their performance isolates the effect of the operators. AIRAgreedy outperforms AIDEgreedy (45.5 % vs. 39.8 %), representing a 14 % relative improvement and underscoring the importance of operators for performance.
Evaluating search policies. Equipped with an improved set of operators, we now explore whether combining them with advanced search methods can further enhance performance. The results (Fig.˜5) show that for R1, AIRAmcts achieves the best performance, achieving state-of-the-art performance on MLE-bench Lite with an average medal rate of 47%. Both AIRAmcts and AIRAgreedy achieve silver-or-above medal rates ($\approx$ 36.5%), outperforming the baseline by 4.5 absolute points. In terms of gold medals, AIRAgreedy performs best, reaching 27%. For o3, AIRAmcts and AIRAgreedy perform comparably in terms of any medal and silver-or-above medal rates, while AIRAmcts outperforms AIRAgreedy in gold medals by 2.2 percentage points. Although agents using R1 and o3 achieve similar overall medal rates, those using o3 perform better when considering gold and silver-or-above medal rates. Specifically, AIRAmcts increases the gold medal rate from 25.68 % with R1 to 30.91 % with o3, a relative increase of 20 %.
Collectively, what stands out is that all search policies combined with AIRA operators outperform AIDEgreedy. Furthermore, the rankings between agents using the AIRA operators with different search strategies differs significantly from that observed in Section˜5.1.
Finally, Fig.˜4 shows the anytime performance of each agent over the 24-hour window defined by MLE-bench. For conciseness, we will focus our discussion on performance with R1, as the results with o3 follow a similar pattern. We observe that the rankings between agents change over time. For example, at the 3-hour mark, the performance gap between AIRAgreedy and AIRAmcts is notable. This gap narrows by the 10-hour mark, where all agents perform similarly. Meaningful divergence in performance appears only after 19 hours. These trends reflect the interaction between rankings and the resources provided, as shown in Figure 7.
We further examine the effect of extended compute time on performance by running experiments for up to 5 days (Appendix 10), showing that our agents continue to improve beyond the 24-hour period. To illustrate differences in search behavior, Appendix 19 presents representative search trees from our experiments for each agent, which may help clarify the strengths and weaknesses of each method.
5.3 The Generalization Gap: Searching with a Proxy Evaluation
[…]
Figure 6: a) The validation–test metric mismatch. Results shown are for agents using R1. Bars depict the absolute performance gap between three configurations: (i) using the validation metric for both intermediate (search) and final (submission) node selection; (ii) using the validation metric only for search and the test metric for selection; and (iii) using the test metric for both search and selection. b) Bridging the validation–test gap. Medal rate achieved by two final node selection strategies as a function of $k$: (i) randomly sampling $k$ nodes and reporting the highest test score among them; (ii) selecting the top $k$ nodes by validation score and reporting the highest test score among those. As $k$ increases, the validation‐based strategy closes the gap to the upper‐bound performance given by the best test score over the entire search graph.
Agents steer their search using the validation scores of candidate solutions, but ultimately, performance is measured on a held-out test set. Due to finite sample effects, the validation score is not perfectly predictive of performance on the test set—a discrepancy known as the generalization gap.
In this section, we measure the impact of this generalization gap on the agents’ performance. We further quantify whether this impact is more detrimental during the intermediate node selection in the search process or the final (submitted) node selection.
How large is the impact of the generalization gap on performance? We begin by comparing two extremes: Val/Val—both intermediate and final node selection use the validation score (standard practice); Test/Test—an oracle baseline using the true test score at both stages. As shown in Fig.˜6, searching based on the test score instead of the validation score improves performance by 9.4 % and 12.4 % for AIRAmcts and AIRAevo. The gap is even higher for the agents using a greedy search policy, reaching 15 % for AIRAgreedy and 16.6 % for AIDEgreedy. For the rest of this section, we investigate the nature of this gap in performance and how to close it.
How much of the gap can be attributed to final node selection? To answer this question, we consider the following two settings: Val/Val; and Val/Test—the intermediate node selection uses the validation score, while the final node selection is made based on the test scores. The results (Fig.˜6) indicate that by selecting the best node in a search graph constructed without privileged access, i.e., based on validation signal, all agents achieve a performance boost of 9 to 13 absolute points. Crucially, comparing Val/Test with Test/Test shows that oracle final-node selection eliminates the gap between the standard (Val/Val) and oracle (Test/Test) settings for both AIRAmcts and AIRAevo. Even for the agents implementing the simpler greedy policy, selecting the best node in the final search graph closes more than 60 % of the gap. These findings highlight robust final-node-selection strategies as a promising path to higher performance.
Bridging the gap through multiple submissions. A straightforward remedy is to hedge against validation score noise by selecting multiple promising nodes. Specifically, among the top-$k$ nodes ranked by validation score in the search graph, we report the test score of the best-performing one. As a baseline, we repeat the process for a set of random-$k$ nodes, and summarize the results in Fig.˜6. We highlight the following observations: 1 the larger gap between top-k and random-k in agents using non-greedy algorithms, AIRAevo and AIRAmcts, suggest greater diversity in their search graphs; 2 with as little as 3 top-k submissions agents achieve an additional 10 % of performance; 3 the top validation nodes are informative of the best performing nodes.
6 Related Works
Scaling Search with LLMs. Integrating search with LLM-based generators has gained traction across domains such as coding (Antoniades et al., 2025), math (Yao et al., 2023a), and planning (Hazra et al., 2024; Hao et al., 2023). In practice, increased test-time compute through algorithms like best-of-NN(Wang et al., 2023), beam search (Hazra et al., 2024), MCTS (Antoniades et al., 2025; Hao et al., 2023), and evolutionary search (Romera-Paredes et al., 2024; Hazra et al., 2025) often improves performance beyond the architectural and size constraints of LLMs (Snell et al., 2024). Our results highlight that these performance gains depend critically on the interplay between search components (see Section˜8)—an aspect that prior work has largely overlooked.
Automating ML Engineering and Scientific Discovery. Traditional approaches like AutoML (Feurer et al., 2022; Thornton et al., 2013; Olson and Moore, 2019) and Neural Architecture Search (NAS)(Zoph and Le, 2017; Liu et al., 2019; Real et al., 2019) automate ML by searching over predefined, expert-designed configuration spaces, often via brute-force or heuristic methods (Bergstra and Bengio, 2012; Elsken et al., 2017; Li et al., 2018). In contrast, recent advances in LLMs enable more open-ended design. AIDE (Jiang et al., 2025) treats ML engineering as LLM-guided code optimization via tree search. AI Scientist v2 (Yamada et al., 2025) extends this paradigm, automating the entire research pipeline using agentic LLMs. Similar techniques have been applied to software engineering (Antoniades et al., 2025), algorithm and reward design (Romera-Paredes et al., 2024; Lu et al., 2024; Faldor et al., 2025; Hazra et al., 2025; DeepSeek-AI, 2025), and even natural sciences (Boiko et al., 2023; Swanson et al., 2024). Existing systems often combine search procedures, operators, and evaluation in ways that make it challenging to understand which components drive performance improvements. To address this, we separate these components and look at how each one—and their interactions—can be optimized better. Concurrent work, R&D-Agent (Yang et al., 2025), which addresses the same problem setting, achieved impressive results on MLE-Bench. Our results fall within their reported standard deviation. We note that their experiments use at most 6 seeds, and as discussed in Section˜16, the limited number of seeds may introduce variation in the conclusions. Finally, our approach differs from methods such as Agent K (Grosnit et al., 2024), which uses long-term memory across multiple Kaggle competitions. In contrast, we focus on addressing each competition independently.
AI Research Frameworks. Most existing benchmarks for machine learning or research engineering, such as MLGym (Nathani et al., 2025), RE-bench (Wijk et al., 2024), and MLAgentBench (Huang et al., 2023), come with their own frameworks. Our approach is most similar to Inspect (Institute, 2025), a framework which provides an abstraction that makes minimal assumptions about the agent and evaluation design, clearly separating the two. This flexibility in agent design is essential for enabling rigorous comparisons across a wide range of agents—from simple LLM-based agents with tool access to scaffolds combining algorithmic and LLM-based components—within the same environment. The environment plays a critical role in performance (see Section˜5.2). However, unlike Inspect, AIRA-dojo focuses strongly on long-running research tasks, which influences our environment design choices. For example, prior work typically uses Docker to containerize agents’ workspaces, but Docker is unsupported on most HPC clusters due to its reliance on root privileges and lack of seamless integration with common HPC resource managers. This limitation hinders scalability, which we address through our Apptainer-based solution (see Section˜3.2).
7 Conclusion
We conceptualize the design of AI research agents as a combination of two axes: search policy and operators. This formulation allowed us to perform a systematic investigation of the interplay between the two, highlighting how the operator set can act as a bottleneck to performance improvements. Based on this finding, we designed an enhanced operator set and constructed agents that combines these operators with several search strategies: Greedy, Monte Carlo Tree Search (MCTS), or Evolutionary Search. Our best-performing agent achieves a new state of the art on MLE-bench, increasing the success rate of winning a Kaggle medal from 39.6 % to 47.7 %. Further, we analyzed the role of the generalization gap in node evaluation. Our findings indicate systematic overfitting: selecting the final solution from the search graph based on its test score instead of its validation score would increase the medal rate by 9 to 13 % (absolute scale), depending on the search strategy. This highlights that robust final-node selection is a promising avenue for improving performance.
7.1 Limitations and Future Work
There are several important dimensions for performance that we leave for future work to explore.
Agentic operators. Our experiments suggest that the effectiveness of search is heavily dependent on the capability of the operators. To manage complexity, in this work we experiment with LLM-based operators. However, one could readily use full-fledged agents as operators. For example, a natural extension would be to include an ideation agent as an operator (Si et al., 2024), and to replace the implementation and debugging operators with a SWE-Agent (Liu et al., 2024; Yang et al., 2024).
Finetuning. LLMs are critical components of the operators. Future work could investigate supervised fine-tuning or reinforcement learning as methods to enhance operator effectiveness.
Scaling the search. To maintain comparability with MLE-bench, we adopt the benchmark’s time (24-hour limit) and compute constraints (1 GPU). However, solving challenging problems likely requires substantially greater resources, and evaluating performance under these restrictions provides limited insights (see Section˜10 for an experiment extended to 120 hours—5 days). Studying the scaling behavior of search policies and operators, and developing agents capable of effectively leveraging more computational resources over longer time horizons, is an important direction for future research.
Data contamination. Finally, there is the issue of data contamination. It is possible that our results are influenced by the presence of information related to the evaluated Kaggle tasks, or similar tasks, within the LLM training data. Creating a continuous stream of fresh and novel tasks remains an important research challenge (White et al., 2025).